Circular No. 7646

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1333 AQUILAE C. Bailyn, S. Tourtellotte, and L. Walkowicz, Yale University; and

D. Gonzalez and J. Espinoza, Cerro Tololo Interamerican Observatory (CTIO), report a new outburst of the optical counterpart of the recurrent x-ray transient Aquila X-1: “R exposures of 900 s made on and before June 13 with the Yale 1-m telescope (+ ANDICAM) at CTIO showed this object at quiescent levels. The R brightness was 0.15 and 0.11 mag brighter than the mean quiescent level on June 14.404 and 15.382 UT, respectively; this is still within 2σ of the quiescent range. However, on June 16.420 the source was 0.36 mag brighter in R than at quiescence, and it has steadily brightened since then, reaching 1.16 mag brighter than quiescence on June 20.361. All magnitudes reported are aperture photometry, and include the near neighbor located 0′′.4 from the true optical counterpart. Previous outbursts have shown x-ray activity starting ∼ 6 days after the start of the optical precursor, so we expect x-ray activity in the very near future.”

SUNGRAZING COMETS Further to IAUC 7642, D. Hammer reports his measurements for six

additional Kreutz sungrazing comets on LASCO C2 coronagraph images obtained with the SOHO spacecraft; comets C/2001 L8 and C/2001 L9 were found by X. Leprette and T. Scarmato, respectively, on images at the SOHO website, while the remaining comets were found by M. Oates. C/1997 K7 appears pointlike with no apparent tail. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2001-M06.

Comet  1997 UT  α_{2000}  δ_{2000}
C/1997 K7 May 31.845  4^{h}32.^{m}4  +20^{o}19′
Comet  2001 UT  α_{2000}  δ_{2000}
C/2001 L5 June 5.814  4^{h}55.^{m}0  +20^{o}53′
C/2001 L6  8.875  5 07.2  +21 09
C/2001 L7  9.004  5 07.6  +21 08
C/2001 L8  10.004  5 10.8  +21 10
C/2001 L9  13.564  5 23.2  +21 34
2001 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green